OpenStudy (anonymous):

I'll give a medal What is the foci of 4x2 + 9y2 = 36

OpenStudy (anonymous):

Write it as \[ \frac {x^2}9 + \frac {y^2}4 =1 \] a=3, b=2; \[ c^2=a^2 - b^2= 9-4=5\\ c=\pm \sqrt 5 Foci=(0,\pm \sqrt 5) \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

i think foci are \[\left( \pm \sqrt{5},0 \right)\]